# Let the Trial Begin: A Mock-Court Approach to Vulnerability Detection using LLM-Based Agents

### v1

@@ 2.4. Review Board (Jury Role)
The review board agent in this VulTrial is similar to the role of jury in a courtroom scenario. In a traditional courtroom, the jury weighs all evidence and testimony presented during a trial, alongside any guidance from the judge, to determine whether a defendant is guilty or not. Similarly, in vulnerability detection, the review board agent is tasked to evaluate the code’s security status based on the evidence presented by both the security researcher and code author agents, as well as the moderator’s summary. Its purpose is to deliver a final verdict on whether the potential vulnerabilities identified by the security researcher are indeed valid. The prompt guiding the review board agent is shown in Figure 6. This agent receives (1) the summary from the moderator agent; (2) the list of potential vulnerabilities from the security researcher agent; (3) the code author agent’s counterarguments; (4) the original code snippet. The review board agent is tasked with generating several outputs.
@@
First, it must extract the names of potential vulnerabilities from the list provided by the security researcher and proceed with a decision for each one. The review board agent needs to highlight the specific vulnerability being discussed as there may be multiple vulnerabilities under review. Second, the agent must determine whether each identified vulnerability is “valid,” “invalid,” or “partially valid.” This process is similar to a jury deciding whether a defendant is guilty or not guilty. Third, the agent must assess the severity of the vulnerability, similar to how a jury determines the extent of damages in a trial. However, instead of assigning an open-ended value, the review board needs to categorize severity as “low,” “medium,” “high.” If the vulnerability is deemed invalid, the severity is marked as “none.” Fourth, the agent must provide a recommended action based on the decision and severity level. For example, if a vulnerability is deemed valid and has a high severity, the recommended action should be “fix immediately.” Finally, the agent must generate an explanation for its decisions. Unlike a jury, which is not required to provide reasoning, we ask the review board agent to do so, as asking LLMs for reasoning can enhance performance (Wei et al., 2022; Kojima et al., 2022; Cheng et al., 2024) and helps users to understand the decision.
@@
We employ a threshold-based approach to classify a function as vulnerable, rather than a simple binary “vulnerable or not,” because it offers flexibility for users to tailor vulnerability detection to their tolerance for risk and false positives. By adjusting the threshold, users can decide how strict they want the classification to be. For instance, if a user prefers to catch even the smallest chance of vulnerability (and is comfortable with a higher rate of false positives), they can opt to flag any partially valid vulnerabilities regardless of severity. Conversely, if a user wants fewer false positives, they can raise the threshold to consider only higher-severity issues as triggers.
@@
In our study, we flag a function as vulnerable if at least one vulnerability raised by the security researcher agent is deemed valid at high severity with recommended action fix immediately by the review board agent. This stricter threshold ensures that only the critical and well-substantiated vulnerabilities contribute to the classification, reducing false positives from lower-risk or ambiguous issues. We aim to reduce noise by excluding vulnerabilities marked as partially valid or those with lower severity levels. Fig. 2 also includes an example of the review board agent’s output. It classifies the first vulnerability (Out-of-Bounds Read) as “valid” with “high” severity. Meanwhile, the second vulnerability (Type Handling) is marked as “partially valid” with “medium” severity. Since at least one confirmed vulnerability is present, the final verdict for this code is vulnerable.
You are the Review Board. After reviewing the Moderator’s summary, Code Author’s and Security Researcher’s argument, and ¡code¿, produce a JSON array of verdicts for each vulnerability identified by the Security Researcher. Each object in the array should include: vulnerability”: The same name as given by the Security Researcher. decision”: One of ‘valid’, ‘invalid’, or ‘partially valid’. severity”: If valid or partially valid, assign a severity (‘low’, ‘medium’, ‘high’); if invalid, use ‘none’. recommended action”: Suggest what should be done next (e.g., ‘fix immediately’, ‘monitor’, ‘no action needed’). reason”: A brief explanation of why you reached this conclusion, considering both the Security Researcher’s and Code Author’s perspectives. ¡code¿
¡chat history¿
Figure 6. Review board agent prompt.
@@ 2.5. Agent Interaction
In VulTrial, four agents interact sequentially: the security researcher agent, the code author agent, the moderator agent, and finally the review board agent. During the first round, the security researcher agent (1) only has access to the original code snippet. In subsequent rounds, it also receives information from the code author and moderator. Similarly, the code author agent (2) initially sees only the security researcher agent’s findings, but in the next rounds it can also incorporate the moderator’s feedback. This setup is similar to a courtroom process, which can allow for multiple rounds of arguments and counterarguments between the prosecutor, defense attorney, and judge. The review board agent (4), mirroring a jury role, provides its verdict only after all discussions have concluded. It does not participate in the debate itself but bases its final decision on the evidence and summaries presented by the other three agents.
Allowing multiple discussion rounds may enhance the iterative refinement of analyses and arguments. It would also provide more chances for the security researcher agent and code author agent for rebuttal. Nevertheless, increased rounds entail higher operational costs. Therefore, it is essential to balance the trade-off between performance improvement and the cost. In this study, we initially limit the number of discussion rounds to one (see Section 4.1), and subsequently, in Section 4.3, we explore how varying the number of discussion rounds impacts the overall performance of VulTrial.
@@ 3.1. Dataset
In our study, we utilized the PrimeVul dataset, recently proposed by Ding et al. (Ding et al., 2025b), to evaluate LLMs using single-agent and multi-agents on vulnerability detection. Specifically, we used the PrimeVul pair data, which presents a challenging benchmark, as the vulnerable and benign code functions share at least 80% of their string content. This dataset ensures no data leakage or exact duplication between the training, testing, and validation sets. The latest version of PrimeVul Pair contains 435 vulnerable and 435 benign functions (Ding et al., 2025a).
@@ 3.2. Research Questions
RQ1: How does the performance of VulTrial compare to that of single-agent and multi-agent baselines on PrimeVul? Prior work (Ding et al., 2025b) demonstrated the poor performance of GPT-3.5 and GPT-4 for vulnerability detection on pairwise vulnerable and benign functions in PrimeVul. These findings highlight the limitations of single-agent prompting strategies in challenging detection scenarios. In this RQ, we investigate whether VulTrial, a multi-agent LLM-based framework (Section 2), can outperform both single-agent methods (Ding et al., 2025b) and existing multi-agent approaches such as GPTLens (Hu et al., 2023) on PrimeVul.
@@
RQ3: Does allowing multiple rounds of debate among role-specific agents in VulTrial lead to improvement compared to a single-pass analysis? In previous RQs, we focus on conducting a single round of discussions among agents to reach a final verdict. However, allowing multiple rounds of discussion could give agents the opportunity to refine their arguments to reach a more accurate decision. In this RQ, we examine whether enabling multiple rounds of discussion can enhance VulTrial’s performance.
RQ4: How does each agent contribute to the overall performance of VulTrial VulTrial features four distinct roles, security researcher, code author, moderator, and review board, which is designed to follow a courtroom-style scenario. While the results from previous RQs highlight VulTrial’s effectiveness, it is not known whether each agent is essential to the final prediction and how it impacts the performance. In this RQ, we do the ablation study that systematically omits one agent at a time to help us verify each agent’s contribution and evaluate whether the loss of a role reduces the performance.
Important note: The *ground-truth label* for this code is: “¡vulnerable/benign¿ vulnerable” means this code contains at least one vulnerability. benign” means there are no vulnerabilities in this code. Use the label to ensure your arguments match the fact: Even if you are “defending” the code, if the label is “vulnerable,” you should not fully deny actual issues. If the label is “benign,” you can confidently refute all vulnerability claims. Here is the CVE ID and description that is related with the code:
¡CVE ID¿”: ¡CVE Description¿
Figure 7. Example of additional information provided to the code author agent for constructing the instruction tuning dataset.
@@ 3.3. Baseline and Models
Baselines. For the single-agent baseline, we utilized the chain of thought (CoT) (Wei et al., 2022) prompting technique, which achieved the highest results in the PrimeVul (Ding et al., 2025b). CoT prompting encourages LLMs to reason through problems step by step to arrive at an answer. For the multi-agent baseline, we utilized an approach from a recent previous study, GPTLens (Hu et al., 2023). GPTLens operates with two types of agents: auditor agent, which identifies potential vulnerabilities, and a critic agent, which evaluates and scores the identified vulnerabilities. The final decision is based on the critic’s score; if the score is less than 5, it is classified as benign; otherwise, it is considered vulnerable. There are several key differences between GPTLens and VulTrial. VulTrial follows a courtroom-inspired setting with four role-specific agents, whereas GPTLens relies on a two-agent structure. Unlike GPTLens, where the critic agent determines the final decision, VulTrial used a different agent (i.e., review board agent) for this task, designed to be neutral toward the security researcher and code author agents. Recent studies (Gao et al., 2024; Hu et al., 2024) suggest that employing domain-specific multi-agent systems that reflect a more realistic scenario leads to better decision making, further supporting our approach in VulTrial.
Models. We employed multiple models in this study. First, we used several well-established OpenAI models known for their strong performance, as confirmed by prior research (Widyasari et al., 2024; Hu et al., 2023; Zhou et al., 2024). Specifically, we utilized GPT-3.5 and GPT-4o.
GPT-3.5 (OpenAI, 2023a): GPT-3.5 is one of the first widely popular OpenAI LLM models. Following PrimeVul (Ding et al., 2025b), we also evaluate the performance of this model in our study. In our evaluation, we utilized gpt-3.5-turbo-0125, the latest iteration of the GPT 3.5 model with a context window of 16k tokens.
GPT-4o (OpenAI, 2024b): We employed GPT-4o, one of OpenAI’s flagship large language models recommended as a replacement for GPT-4 (OpenAI, 2023b). We chose GPT-4o over GPT-4 (used in PrimeVul’s evaluation) because OpenAI indicates that GPT-4o offers performance comparable to GPT-4 while significantly reducing computational costs (OpenAI, 2024b). Specifically, we used gpt-4o-2024-08-06 allows up to 128k tokens in its context window and 16k output tokens.
@@ 3.4. Evaluation Metrics
Following the approach in PrimeVul (Ding et al., 2025b), we adopt a pair-wise evaluation strategy. It is designed to assess the model’s prediction on pairs of vulnerable and benign code, rather than the single codes. This method highlights the model’s capacity to differentiate subtle variations that make one function vulnerable and the other benign. The four key metrics are defined as follows:
  • Pair-wise Correct Prediction (P-C): The model correctly predicted both codes in the pair.
  • Pair-wise Vulnerable Prediction (P-V): The model incorrectly classifies both code as vulnerable.
  • Pair-wise Benign Prediction (P-B): The model incorrectly classifies both code as benign.
  • Pair-wise Reversed Prediction (P-R): The model assigns the opposite label to each code in the pair, predicting “benign” for a vulnerable function and “vulnerable” for a benign one.
We use P-C as our primary evaluation metric. A high P-C indicates accurate vulnerability detection that avoids false positives and negatives, making it a more robust indicator of overall performance.
@@ 3.5.1. Answering RQ1:
We compare VulTrial with both single-agent (Ding et al.’s CoT(Ding et al., 2025b)) and multi-agent (GPTLens) baselines, using two different LLMs, i.e., GPT-3.5 and GPT-4o. Section 3.3 provides further details on our baseline methods and model choices. We evaluate each approach using several metrics, with the Pair-wise Correct (PC) rate as our primary evaluation metric. A higher PC rate indicates better performance.
@@ 3.5.2. Answering RQ2:
For this RQ, we conducted instruction tuning using samples drawn from GPT-4o’s validation set. After running GPT-4o on the validation set, we took its correct outputs to create an instruction tuning dataset for a GPT-4o model. We selected GPT-4o because, as shown in RQ1, VulTrial using GPT-4o outperforms both single-agent and multi-agent baselines and achieves the best overall results, making it an ideal candidate for evaluating the impact of instruction tuning. Our goal is to determine whether instruction tuning can further enhance its performance. We chose GPT-4o for generating the instruction data to ensure consistency in style and reasoning patterns. By using the same models for generating data, we reduce the risk of mismatches that can occur when instruction tuning with outputs from a completely different model.
@@
For instruction tuning data, we use 50 code pairs (vulnerable and benign) with the fewest tokens in the validation set to help manage the tuning cost. In addition, OpenAI has highlighted that GPT-4o can achieve strong results with “as few as a few dozen examples” (OpenAI, 2024a). When creating our instruction tuning dataset, we initially included the true label in the prompt (e.g., adding vulnerability-related information for the code author agent, see Fig. 7), but then removed direct label disclosures in the final dataset. As a result, each prompt retained its original role-specific structure and was paired with the correct assistant output.
@@
We utilized following scenarios for the evaluation: (1) no instruction tuning, (2) instruction tuning only the security researcher agent, (3) instruction tuning only the code author agent, (4) instruction tuning only the moderator agent, (5) instruction tuning only the review board agent, (6) instruction tuning both security researcher and code author agents, (7) instruction tuning the security researcher, code author, and moderator agents, and (8) instruction tuning all agents. To systematically assess the impact of instruction tuning, we first applied it to individual agents to understand their specific contributions. Then, we incrementally introduced instruction tuning in the order of agent interaction, allowing us to analyze how tuning at different stages influences the overall performance of VulTrial. This stepwise approach provides insights into the cumulative effects of instruction tuning across multi-agents.

### v2

@@ 2.4. Review Board (Jury Role)
In a traditional courtroom, the jury weighs all evidence and testimony presented during a trial, alongside any guidance from the judge, to determine whether a defendant is guilty or not. Similarly, in vulnerability detection, the review board agent is tasked to evaluate the code’s security status based on the evidence presented by both the security researcher and code author agents, as well as the moderator’s summary. Its purpose is to deliver a final verdict on whether the potential vulnerabilities identified by the security researcher are indeed valid. This agent receives (1) the summary from the moderator agent; (2) the list of potential vulnerabilities from the security researcher agent; (3) the code author agent’s counterarguments; (4) the original code snippet.
@@
The review board agent is tasked with generating several outputs. First, it must extract the names of potential vulnerabilities from the list provided by the security researcher and proceed with a decision for each one. Second, the agent must determine whether each identified vulnerability is “valid,” “invalid,” or “partially valid.” This process is similar to a jury deciding whether a defendant is guilty or not guilty. Third, the agent must assess the severity of the vulnerability, similar to how a jury determines the extent of damages in a trial. However, instead of assigning an open-ended value, the review board needs to categorize severity as “low,” “medium,” or “high.” If the vulnerability is deemed invalid, the severity is marked as “none.” Fourth, the agent must provide a recommended action based on the decision and severity level. For example, if a vulnerability is deemed valid and has a high severity, the recommended action should be “fix immediately.” Finally, the agent must generate an explanation for its decisions. Unlike a real jury, we require the review board agent to provide reasoning, as it improves LLMs’ performance (wei2022chain; kojima2022large; cheng2024chainlm) and helps users understand the decision.
@@
We use a threshold rather than a simple binary “vulnerable or not,” because it offers flexibility for users to tailor vulnerability detection to their tolerance for risk and false positives. For instance, if a user prefers to catch even the smallest chance of vulnerability, they can opt to flag any partially valid vulnerabilities. Conversely, if a user wants fewer false positives, they can raise the threshold to consider only higher-severity issues.
@@
We flagged a function as vulnerable if at least one vulnerability raised by the security researcher agent is deemed valid at high severity with recommended action fix immediately by the review board agent. This threshold ensures that only the critical and well-substantiated vulnerabilities contribute to the classification, reducing false positives from lower-risk or ambiguous issues. We aim to reduce noise by excluding vulnerabilities marked as partially valid or those with lower severity levels. Fig. 2 also includes an example of the review board agent’s output. It classifies the first vulnerability (Out-of-Bounds Read) as “valid” with “high” severity. Meanwhile, the second vulnerability (Type Handling) is marked as “partially valid” with “medium” severity. As at least one confirmed vulnerability is present, the final verdict for this code is vulnerable.
@@ 2.5. Agent Interaction
In VulTrial, four agents interact sequentially: the security researcher agent, the code author agent, the moderator agent, and finally the review board agent. During the first round, the security researcher agent (1) has access only to the original code snippet. In subsequent rounds, it also receives information from the code author and moderator. Similarly, the code author agent (2) initially sees only the security researcher agent’s findings, but in the next rounds, it can also incorporate the moderator’s (3) output. This setup is similar to a courtroom process, which can allow for multiple rounds of arguments and counterarguments between the prosecutor, defense attorney, and judge. The review board agent (4), mirroring a jury role, provides its verdict only after all discussions have concluded. It does not participate in the debate itself but bases its final decision on the evidence and summaries presented by the other three agents.
Allowing multiple discussion rounds may enhance the iterative refinement of analyses and arguments. It also provides both the security researcher and code author agents with more opportunities for rebuttal. In this study, we initially limit the number of discussion rounds to one (Section 4.1) and then assess how changing the number of rounds affects VulTrial’s performance (Section 4.3).
@@ 3.1. Dataset
We utilized the PrimeVul dataset, recently proposed by Ding et al. (ding2024vulnerability), to evaluate vulnerability detection approaches. Specifically, we used the PrimeVul pair dataset, which presents a challenging benchmark, as the vulnerable and benign code functions share at least 80% of their string content. The pair dataset includes 3,789 training pairs, 480 validation pairs, and 435 test pairs, with no data leakage across splits. In contrast, the full PrimeVul dataset contains 175,797 training functions (4,862 vulnerable and 170,935 benign) and 23,948 validation functions (593 vulnerable and 23,355 benign).
@@ 3.2. Research Questions
RQ1: How does the performance of VulTrial compare to that of single-agent and multi-agent baselines on PrimeVul? Prior work (ding2024vulnerability) demonstrated the poor performance of GPT-3.5 and GPT-4 for vulnerability detection on pairwise vulnerable and benign functions in PrimeVul. These findings highlight the limitations of single-agent prompting strategies for this challenging setting. In this RQ, we investigate whether VulTrial, a multi-agent LLM-based framework (Section 2), can outperform both single-agent methods (ding2024vulnerability) and existing multi-agent approaches such as GPTLens (hu2023large) on the PrimeVul pair.
@@
RQ3: Does allowing multiple rounds of debate among role-specific agents in VulTrial lead to improvement compared to a single-pass analysis? In previous RQs, we focused on conducting a single round of discussions among agents to reach a final verdict. However, allowing multiple rounds of discussion could give agents the opportunity to refine their arguments to reach a more accurate decision. In this RQ, we examine whether enabling multiple rounds of discussion can enhance VulTrial’s performance.
RQ4: How does each agent contribute to the overall performance of VulTrial VulTrial features four distinct roles, security researcher, code author, moderator, and review board, which is designed to follow a courtroom-style scenario. While the results from previous RQs highlight VulTrial’s effectiveness, it is not known whether each agent is essential to the final prediction and how it impacts the performance. In this RQ, we do the ablation study that omits one agent at a time to verify each agent’s contribution.
@@ 3.3. Baseline and Models
Baselines. We employed both DL-based and LLM-based vulnerability detection techniques as single-agent baselines. For the DL-based methods, we used several state-of-the-art models widely adopted in prior vulnerability detection studies (ding2024vulnerability; xiong2024vuld; zhou2024large; chen2023diversevul):
  • CodeBERT (feng2020codebert): A bimodal pre-trained transformer model trained on source code and natural language.
  • CodeT5 (wang2021codet5): A transformer model that learns to understand and generate code across multiple programming languages.
  • UniXCoder (guo2022unixcoder): A cross-modal pre-trained model that jointly learns representations of code, ASTs, and execution traces.
  • LineVul (fu2022linevul): A line-level transformer-based vulnerability prediction model detecting vulnerable lines through fine-grained context representations.
For CodeBERT, CodeT5, and UniXCoder, we performed two separate fine-tuning: (1) using the full PrimeVul training and validation sets, and (2) using only the pair-wise subsets from PrimeVul. For LineVul, we first evaluated the model provided in the previous study (fu2022linevul). In addition, we further fine-tuned LineVul using the pair-wise subset to ensure a consistent comparison with the other DL models. For the LLM-based approach, we utilized the chain of thought (CoT) (wei2022chain) prompting technique, which achieved the highest results in the PrimeVul (ding2024vulnerability). CoT prompting encourages LLMs to reason through problems step by step to arrive at an answer.
For the multi-agent baseline, we utilized an approach from a recent study, GPTLens (hu2023large). GPTLens operates with two types of agents: auditor agent, which identifies potential vulnerabilities, and a critic agent, which evaluates and scores the identified vulnerabilities. The final decision is based on the critic’s score; if the score is less than 5, it is classified as benign; otherwise, it is considered vulnerable. There are several key differences between GPTLens and VulTrial. VulTrial follows a courtroom-inspired setting with four role-specific agents, whereas GPTLens relies on a two-agent structure. Unlike GPTLens, where the critic agent determines the final decision, VulTrial used a different agent (i.e., review board agent) for this task, designed to be neutral toward the security researcher and code author agents. Recent studies (gao2024large; hu2024unlocking) suggest that employing domain-specific multi-agent systems that reflect a more realistic scenario leads to better decision making, further supporting our approach in VulTrial.
Models. We used well-established OpenAI models known for their strong performance, as confirmed by prior research (widyasari2024beyond; hu2023large; zhou2024large_emerging).
GPT-3.5 (openai2023gpt35): GPT-3.5 is one of the first widely popular OpenAI LLM models. Following PrimeVul (ding2024vulnerability), we also evaluate the performance of this model in our study. In our evaluation, we utilized gpt-3.5-turbo-0125, the latest iteration of the GPT 3.5 model with a context window of 16k tokens.
GPT-4o (openai2024gpt4o): We chose GPT-4o, one of OpenAI’s flagship models, over GPT-4 (openai2023gpt4) (used in PrimeVul’s evaluation) because OpenAI indicates that GPT-4o offers performance comparable to GPT-4 while significantly reducing computational costs (openai2024gpt4o). Specifically, we used gpt-4o-2024-08-06 allows up to 128k tokens in its context window and 16k output tokens.
@@ 3.4. Evaluation Metrics
Following the approach in PrimeVul (ding2024vulnerability), we adopt a pair-wise evaluation strategy. It is designed to assess the model’s prediction on pairs of vulnerable and benign code, rather than the single codes. This method highlights the model’s capacity to differentiate subtle variations that make one function vulnerable and the other benign. We report five metrics. Pair-wise Correct Prediction (P-C) records pairs where both functions are classified correctly. Pair-wise Vulnerable Prediction (P-V) records pairs where both functions are incorrectly labeled as vulnerable, while Pair-wise Benign Prediction (P-B) records pairs labeled incorrectly as benign. Pair-wise Reversed Prediction (P-R) captures cases where the model assigns opposite labels to the pair, predicting “benign” for the “vulnerable” code and “vulnerable” for the “benign” one. Pair-wise Error (E) reflects pairs in which one function is misclassified. Among these metrics, P-C is our primary evaluation metric. A high P-C indicates accurate vulnerability detection that avoids false positives and negatives, making it a more robust indicator of overall performance.
To provide more comprehensive discussion in VulTrial and baselines comparison, we also report standard classification metrics: Precision (P), Recall (R), and False Positive Rate (FPR), following the formula used in the previous study (ding2024vulnerability). These traditional metrics provide additional insights into the model’s general classification behavior, particularly in individual predictions.
@@ 3.5.1. Answering RQ1:
We compare VulTrial with both single-agent (CodeBERT (feng2020codebert), CodeT5 (wang2021codet5), UniXCoder (guo2022unixcoder), LineVul (fu2022linevul), and Ding et al.’s CoT(ding2024vulnerability)) and multi-agent (GPTLens) baselines. For the LLM-based techniques, we utilize two different LLMs, i.e., GPT-3.5 and GPT-4o (see Section 3.3). We evaluate each approach using several metrics (P-C, P-V, P-B, P-R, Error, P, R, and FPR).
@@ 3.5.2. Answering RQ2:
For this RQ, we conducted instruction tuning using samples drawn from GPT-4o’s validation set. After running GPT-4o on the set, we collected its correct outputs to build an instruction-tuning dataset for a GPT-4o model. We focus on GPT-4o because, as shown in RQ1, VulTrial with GPT-4o outperforms both single-agent and multi-agent baselines and achieves the best overall results. Our goal is to determine whether instruction tuning can further enhance its performance. We chose GPT-4o for generating the instruction data to ensure consistency in style and reasoning patterns. By using the same models for generating data, we reduce the risk of mismatches that can occur when instruction tuning with outputs from a completely different model.
@@
For instruction tuning data, we use 50 code pairs (vulnerable and benign) with the fewest tokens in the validation set to help manage the tuning cost. We chose 50 samples based on preliminary experiments comparing different dataset sizes (25, 50, and 75 pairs), as 50 resulted in significant performance gains over 25 samples, while achieving comparable results to using 75 samples (ourreplicationpackageall). In addition, OpenAI has highlighted that GPT-4o can achieve strong results with “as few as a few dozen examples” (openai2024gpt4o_finetuning). When creating our instruction tuning dataset, we initially included the true label in the prompt, but then removed direct label disclosures in the final dataset. For example, for the code author agent (ourreplicationpackageall), we included vulnerability-related information such as CVE ID and label. Similarly, for the security researcher agent, we added the CVE description and label for both vulnerable and benign code, prompting the agent to connect code to relevant vulnerabilities. For the review board agent, the prompt included ground-truth labels but also emphasized that the agent should not reference them explicitly in its reasoning. In contrast, the moderator agent’s prompt excluded any ground-truth vulnerability information and was cleaned to avoid indirect label leakage from chat history. As a result, each prompt retained its original role-specific structure and was paired with the correct assistant output.
For this instruction tuning data, we ensure that the agent’s predicted label (“vulnerable” or “benign”) matches the ground-truth label. Additionally, we manually verify the quality of the agent’s explanation, confirming that it correctly identifies the presence or absence of a vulnerability. An example of the vulnerable instance is shown in the online appendix (ourreplicationpackageall); this vulnerability is classified as CWE-416: Use After Free. The security researcher agent correctly identifies the issue within the CleanWriters function, where the sequence of operations may lead to undefined behavior if gf_isom_box_del() uses the writer object after it has been freed. The explanation aligns well with the root cause described in CVE-2020-35980. For the corresponding benign variant (ourreplicationpackageall), the code author agent correctly refutes the presence of a vulnerability by noting that gf_list_del_item() only removes the items from the list without deallocating their memory. The justification provided by the review board agent also matches the code author’s agent clarification and supports the benign ground-truth label.
@@
We evaluated eight scenarios: (1) no instruction tuning; (2-5) tuning only the security researcher, code author, moderator, or review board agent; (6) tuning security researcher and code author; (7) tuning security researcher, code author, and moderator; and (8) tuning all agents. To systematically assess the impact of instruction tuning, we first applied it to individual agents to understand their specific contributions. Then, we incrementally introduced instruction tuning in the order of agent interaction, allowing us to analyze how tuning at different stages influences the overall performance of VulTrial. This stepwise approach provides insights into the cumulative effects of instruction tuning across multiple agents.